
N10 amplifier issues returned in the second-half shift; exposures 775154-775157 affected. Chip reset and looked good for 2 hours.
N10 amplifier issue returned for a second time in the same half-shift. Affected exposures 775189-775204. Chip was reset, bias frame looked correct, but wide-field exposures 775206-775208 afterwards had same issue. Reset again and standards looked good.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2018-09-17 | 10.0 | 0.0 | 0.0 | 0.0 | 0.0 | 10 |
Links
[1] mailto:djschlegel@lbl.gov
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7251/delete?destination=print/node/14546
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7251/edit?destination=print/node/14546
[4] mailto:frieman@fnal.gov
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7252/delete?destination=print/node/14546
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7252/edit?destination=print/node/14546
[7] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/14546?destination=print/node/14546
DES nightsum: http://des-ops.fnal.gov:8080/nightsum/nightsum-2018-09-17/nightsum.html